# Spin Inertia as a Source of Topological Magnons: Chiral Edge States from Coupled Precession and Nutation

### 06 Mar 2026 v1

@@ Abstract
The topology of the band structures of spin waves or magnons has also attracted considerable attention recently [20]. Topological magnon insulators with chiral edge modes have been proposed theoretically in multi-sublattice ferromagnets [21, 22, 23, 24], magnon crystals [25] and skyrmion lattices [26, 27, 28], and magnonic Weyl nodes [29, 30] have also been identified. The topology of the band structure also contributes to transport phenomena such as the magnon thermal Hall effect [31, 32, 33, 34]. A prototypical example of a topological magnon insulator is a ferromagnet on a honeycomb lattice, realized in, e.g., chromium trihalides and Cr(Si/Ge)Te3 compounds [35, 36, 37]. Two different microscopic mechanisms have been put forward that could explain the topologically non-trivial nature of the band structure in this structure: the Dzyaloshinsky–Moriya interaction [38, 39, 35, 36] which is antisymmetric in the lattice sites, and various forms of the symmetric two-site exchange anisotropy including dipolar, pseudodipolar, and Kitaev terms [40, 41, 42]. It has been suggested that the Dzyaloshinsky–Moriya interaction may be more prominent in the opening of a gap in the band structure [43], since the gap size scales linearly with its strength but quadratically with the pseudodipolar term. Due to their different symmetry properties, it was also proposed that the two contributions may be distinguished by measuring the magnon spectrum for different orientations of the magnetization [44].
Here, we present how spin inertia influences the topology of the magnon band structure. As an example, we demonstrate that the precessional and nutational magnon bands may hybridize in a honeycomb ferromagnet, leading to an additional gap opening in the spectrum. This hybridization is enabled by the pseudodipolar interaction but not by the Dzyaloshinsky–Moriya interaction, enabling to distinguish between these two contributions, and the gap size is shown to scale linearly with the interaction strength. We confirm the topological nature of this gap by demonstrating the emergence of edge modes in slab geometries, and connect this to the Chern numbers of the bands.
@@
Here, $\bm{S}_{i}$ denotes a unit vector at site $i$ corresponding to the direction of the magnetic moment, $\gamma$ is the gyromagnetic ratio, $\alpha$ represents the Gilbert damping constant, and $\eta$ is the inertial relaxation time. The effective magnetic field is expressed as $\bm{B}^{\rm eff}_{i}=-\frac{1}{M_{i}}\frac{\partial\mathcal{H}}{\partial\bm{S}_{i}}$, where $\mathcal{H}$ is the Hamiltonian of the system and $M_{i}$ is the magnetic moment. The angular momentum at site $i$ may be defined as $\bm{L}_{i}=\frac{M_{i}}{\gamma}\left(\bm{S}_{i}-\eta\bm{S}_{i}\times\dot{\bm{S}}_{i}\right)$ [8], see the Supplemental Material [45]. In the absence of an external magnetic field and spin–orbit coupling, the system is invariant under global rotations, which leads to the conservation of the total angular momentum $\sum_{i}\bm{L}_{i}$.
@@
If the calculations are performed around a stable equilibrium state, $\mathcal{H}_{\textrm{SW}}$ is a positive definite operator, and the eigenvalue problem is quasi-Hermitian [46]. This means that all eigenvalues are real, but the left $\left<\bm{l}_{q}\right|$ and right $\left|\bm{r}_{q}\right$> eigenvectors belonging to the same eigenvalue are not adjoints of each other, but form a biorthogonal system instead, $\left<\bm{l}_{q}|\bm{r}_{q^{\prime}}\right>=\delta_{qq^{\prime}}$. Furthermore, the eigenvalues are found in $\pm\omega_{q}$ pairs due to the particle-hole symmetry of the system, which is represented as $\mathcal{C}=\textrm{diag}\left(\bm{\sigma}^{x},-\bm{\sigma}^{x}\right)\mathcal{K}$ in this system, where $\mathcal{K}$ denotes complex conjugation. All eigenvectors have the form $\left|\bm{r}_{q}\right>=\left[\bm{v}_{q},\omega_{q}\bm{v}_{q}\right]$, since the latter components are connected to the spin velocity.
In the following, we consider a ferromagnetic system magnetized along the global $z$ direction, and describe the undamped modes for $\alpha=0$. The rotational symmetry around the $z$ axis is preserved in the absence of spin–orbit coupling, and the $z$ component of the total angular momentum, $\sum_{i}L^{z}_{i}$, remains a conserved quantity. On the level of spin-wave theory, this is reflected by the absence of pairing terms in the Hamiltonian $\mathcal{H}_{\textrm{SW}}$, which would couple the first and second components of $\bm{S}^{\bot}$ together. Considering only the $S_{i}^{+}$ sector, two types of excitations may be distinguished in inertial dynamics [47]: precessional excitations performing a counterclockwise rotation around the direction of the total angular momentum, which have $\omega>0$; and nutational excitations performing a clockwise rotation with $\omega<0$. These are illustrated in Fig. 1. Due to the different rotational senses, the angular momenta of the excited states also differ: $\sum_{i}L^{z}_{i}>\frac{M_{i}}{\gamma}\sum_{i}S^{z}_{i}$ for precessional and $\sum_{i}L^{z}_{i}<\frac{M_{i}}{\gamma}\sum_{i}S^{z}_{i}$ for nutational motion. To observe topological edge modes connecting the precessional and nutational magnon bands, this conservation of the total angular momentum has to be broken.
Figure 1: Sketch of the ferromagnet on the honeycomb lattice. Red and blue trajectories denote counterclockwise precessional and clockwise nutational modes, respectively.
@@
where $A$ and $B$ indicate the two sublattices denoted by red and blue spins in Fig. 1, respectively. The exchange interaction tensor $\bm{J}_{ij}$ between nearest neighbors is diagonal here, with $J^{zz}<J^{xx}=J^{yy}<0$ preferring magnetization along $\hat{\bm{z}}$. The pseudodipolar interaction $F$ connects the spin directions to the relative positions of the sites via the unit vector $\hat{\bm{r}}_{ij}$ connecting sites $i$ and $j$, which is either $\hat{{\bm{\delta}}}_{1}=\left(-1,0\right)$, $\hat{\bm{\delta}}_{2}=\left(\frac{1}{2},-\frac{\sqrt{3}}{2}\right)$, or $\hat{\bm{\delta}}_{3}=\left(\frac{1}{2},\frac{\sqrt{3}}{2}\right)$. The parameter values used in the simulations are summarized in Table 1.
@@
$\mathcal{H}_{\textrm{SW}}=$ $\begin{pmatrix}\Omega&\mathcal{H}_{1}&0&\tilde{F}_{1}-{\rm i}\tilde{F}_{2}\\[10.0pt] \mathcal{H}_{1}^{*}&\Omega&\tilde{F}_{1}^{*}-{\rm i}\tilde{F}^{*}_{2}&0\\[10.0pt] 0&\tilde{F}_{1}+{\rm i}\tilde{F}_{2}&\Omega&\mathcal{H}_{1}\\[10.0pt] \tilde{F}_{1}^{*}+{\rm i}\tilde{F}^{*}_{2}&0&\mathcal{H}_{1}^{*}&\Omega\end{pmatrix}\,$, (5)
where the components are given in the order $\left[S_{-}^{A},S_{-}^{B},S_{+}^{A},S_{+}^{B}\right]$. The matrix elements are given by
@@
$\mathcal{H}_{1}\left(\bm{k}\right)=$ $\left(J^{xx}+\frac{F}{2}\right)\left(\textrm{e}^{\textrm{i}\bm{k}\delta_{1}}+\textrm{e}^{\textrm{i}\bm{k}\delta_{2}}+\textrm{e}^{\textrm{i}\bm{k}\delta_{3}}\right),$ (7)
$\tilde{F}_{1}\left(\bm{k}\right)=$ $\frac{F}{2}\left(2\textrm{e}^{\textrm{i}\bm{k}\delta_{1}}-\textrm{e}^{\textrm{i}\bm{k}\delta_{2}}-\textrm{e}^{\textrm{i}\bm{k}\delta_{3}}\right),$ (8)
$\tilde{F}_{2}\left(\bm{k}\right)=$ $\frac{\sqrt{3}F}{4}\left(\textrm{e}^{\textrm{i}\bm{k}\delta_{3}}-\textrm{e}^{\textrm{i}\bm{k}\delta_{2}}\right).$ (9)
The calculated magnon spectrum is shown in Fig. 2 along the high-symmetry lines of the Brillouin zone. The exchange interaction forms two precessional bands in this two-sublattice structure, which touch in the $K$ and $K^{\prime}$ points forming Dirac cones. Spin inertia leads to additional nutational modes, which have the same dispersion apart from being shifted up in frequency by $\eta^{-1}$ [17]; see the Supplemental Material [45]. For sufficiently high values of $\eta\gamma M^{-1}\left|J^{xx}\right|$, first the top of the higher precession band touches the bottom of the lower nutation band in the $\Gamma$ point, then a nodal line is formed between the bands along a contour encompassing the $\Gamma$ point. The pseudodipolar interaction $F$ breaks the continuous rotational symmetry around $\hat{\bm{z}}$, thereby causing a hybridization between the precessional and nutational modes and opening gaps in the spectrum. In the Dirac points, the gap is opened for any finite value of $F$, but its size scales with $\left|F\right|^{2}$, making it difficult to identify in Fig. 2. Along the nodal line, the gap size scales linearly with $\left|F\right|$, but this effect can only be observed if the inertial relaxation time is sufficiently high such that the bands overlap. The hybridization between the nutational and precessional bands may be captured by calculating the ellipticity of mode $n$ at wave vector $\bm{k}$,
@@
Figure 3: Magnon band structure in a slab geometry. The system is 4040 unit cells wide along the $x$ direction with open boundary conditions, while it is infinite along the $y$ direction with periodic boundary conditions, meaning that the eigenstates can be indexed by the wave vector component $k_{y}$. $a$ denotes the distance between nearest-neighbor atoms in the lattice, i.e., the lattice constant is $\sqrt{3}a$. Dotted lines show the boundaries of the one-dimensional Brillouin zone. Coloring shows the localization defined in Eq. (11).
To confirm the topologically non-trivial origin of the edge states, we calculate the topological invariants of the bands in the infinite system, corresponding to the Chern number in the considered case. The Chern number is the integral of the $z$ component of the Berry curvature $C_{n}=\left(2\pi\right)^{-1}\int B^{z}_{n}\left(\bm{k}\right)\textrm{d}^{2}\bm{k}$, where the latter may be expressed with the left and right eigenvectors as (see the Supplemental Material [45] for details)
@@
Here, $\mathcal{D}_{\bm{k}}$ is the matrix on the right-hand side of Eq. (2) in Fourier space, and $\alpha,\beta,\gamma$ denote Cartesian indices. The distribution of $B^{z}_{n}$ for the four bands is shown in Fig. 4. The sum of the Chern numbers $\left|\sum_{i=1}^{n}C_{n}\right|$ predicts the number of pair of chiral edge modes between bands $n$ and $n+1$, which equals $1$ between any two bands, in agreement with Fig. 3. The Chern number $C_{1}=-1$ of the lowest band stems in Fig. 4(a) from the peaks in the Berry curvature close to the $K$ and $K^{\prime}$ points where the degeneracy is lifted. These peaks are also observable in the second band in Fig. 4(b) with opposite sign, but they are cancelled out by the circular depression around the contour where the precessional and nutational bands overlap in the absence of spin–orbit coupling. The Berry curvatures in the third and fourth bands are qualitatively similar to the second and first bands, respectively, although with reversed signs.
@@
In summary, we demonstrated that the hybridization between nutational and precessional spin waves in inertial spin dynamics may modify the topological properties of the bands. We illustrated this on the example of a honeycomb ferromagnet, which realizes a magnonic Chern insulator with an avoided crossing between the precessional and nutational bands in the bulk system, which is filled by chiral edge modes in a slab geometry. Assuming an exchange interaction strength of $J\approx-3$ meV and a magnetic moment of $M=3\penalty 10000\ \mu_{\textrm{B}}$, which are in the range of experimentally determined values for honeycomb van der Waals materials [35, 48, 36, 37], the inertial relaxation time in the simulations was around $\eta\approx 300$ fs, which is similar to its value determined in transition metals [10, 8]. Note that the value of $|J|$ may be higher in these materials than what was determined from a non-inertial model, since inertial effects considerably reduce the width of the magnon bands [17]; and that the inertial relaxation time may be further enhanced in the presence of heavy elements with high spin–orbit coupling. Therefore, this phenomenon might be observable in experiments.
These effects may be probed by measuring the magnon spectrum in the whole Brillouin zone up to higher energies where the nutational bands are expected to emerge. The nutational and precessional bands may be distinguished based on their different angular momentum, or opposite rotational sense. This means that methods sensitive to the rotational sense of the excitations may be particularly suitable for this purpose, such as spin-polarized electron-energy loss spectroscopy. Optical methods may also be applied in the vicinity of the $\Gamma$ point where the reversed rotational sense of the higher modes already provides an indication for the topologically non-trivial nature of the bands, as discussed above. Detecting the inertial dynamics may also shed further light on the role of the competing mechanisms of the Dzyaloshinsky–Moriya interaction and the pseudodipolar interaction in the gap opening in these systems [36, 44]: while the pseudodipolar interaction may open a gap between the precessional and nutational bands, the out-of-plane component of the Dzyaloshinsky–Moriya interaction has no such effect since it preserves the angular momenta of the excitations.
These calculations may be generalized to other systems, for example single-sublattice ferromagnets, where topologically non-trivial bands may only be formed if inertial effects are taken into account; see the Supplemental Material [45] for an example. The generation of bulk and topological edge modes in ferromagnets at THz frequencies may also open further perspectives in controlling the angular-momentum transfer between magnons and phonons or electrons.

### 31 Mar 2026 v2

@@ Abstract
The topology of the band structure of spin waves or magnons has also attracted considerable attention recently [20]. Topological magnon insulators with chiral edge modes have been proposed theoretically in multi-sublattice ferromagnets [21, 22, 23, 24], magnon crystals [25] and skyrmion lattices [26, 27, 28], and magnonic Weyl nodes [29, 30] have also been identified. The topology of the band structure also contributes to transport phenomena such as the magnon thermal Hall effect [31, 32, 33, 34]. A prototypical example of a topological magnon insulator is a ferromagnet on a honeycomb lattice, realized in, e.g., chromium trihalides and Cr(Si/Ge)Te3 compounds [35, 36, 37]. Two different microscopic mechanisms have been put forward that could explain the topologically non-trivial nature of the band structure in this geometry: the Dzyaloshinsky–Moriya interaction [38, 39, 35, 36] which is antisymmetric in the lattice sites, and various forms of the symmetric two-site exchange anisotropy including dipolar, pseudodipolar, and Kitaev terms [40, 41, 42]. It has been suggested that the Dzyaloshinsky–Moriya interaction may be more prominent in the opening of a gap in the band structure [43], since the gap size scales linearly with its strength but quadratically with the pseudodipolar term. Due to their different symmetry properties, it was also proposed that the two contributions may be distinguished by measuring the magnon spectrum for different orientations of the magnetization [44].
Here, we present how spin inertia influences the topology of the magnon band structure. As an example, we demonstrate that the precessional and nutational magnon bands may hybridize in a honeycomb ferromagnet, leading to an additional gap opening in the spectrum. This hybridization is enabled by the pseudodipolar interaction but not by the Dzyaloshinsky–Moriya interaction, enabling to distinguish between these two contributions, and the gap size is shown to scale linearly with the interaction strength. We confirm the topological nature of this gap by demonstrating the emergence of chiral edge modes in slab geometries, and connect this to the Chern numbers of the bands.
@@
Here, $\bm{S}_{i}$ denotes a unit vector at site $i$ corresponding to the direction of the magnetic moment, $\gamma$ is the gyromagnetic ratio, $\alpha$ represents the Gilbert damping constant, and $\eta$ is the inertial relaxation time. The effective magnetic field is expressed as $\bm{B}^{\rm eff}_{i}=-\frac{1}{M_{i}}\frac{\partial\mathcal{H}}{\partial\bm{S}_{i}}$, where $\mathcal{H}$ is the Hamiltonian of the system and $M_{i}$ is the magnetic moment. The angular momentum at site $i$ may be defined as $\bm{L}_{i}=\frac{M_{i}}{\gamma}\left(\bm{S}_{i}-\eta\bm{S}_{i}\times\dot{\bm{S}}_{i}\right)$ [8], see the Supplemental Material (SM) [45]. In the absence of an external magnetic field and spin–orbit coupling, the system is invariant under global rotations, which leads to the conservation of the total angular momentum $\sum_{i}\bm{L}_{i}$.
@@
If the calculations are performed around a stable equilibrium state, $\mathcal{H}_{\textrm{SW}}$ is a positive definite operator, and the eigenvalue problem is quasi-Hermitian [46]. This means that all eigenvalues are real, but the left $\left<\bm{l}_{q}\right|$ and right $\left|\bm{r}_{q}\right$> eigenvectors belonging to the same eigenvalue are not adjoints of each other, but form a biorthogonal system instead, $\left<\bm{l}_{q}|\bm{r}_{q^{\prime}}\right>=\delta_{qq^{\prime}}$. Furthermore, the eigenvalues are found in $\pm\omega_{q}$ pairs due to the particle-hole symmetry, which is represented as $\mathcal{C}=\textrm{diag}\left(\bm{\sigma}^{x},-\bm{\sigma}^{x}\right)\mathcal{K}$ in this system, where $\mathcal{K}$ denotes complex conjugation. All eigenvectors have the form $\left|\bm{r}_{q}\right>=\left[\bm{v}_{q},\omega_{q}\bm{v}_{q}\right]$, since the latter components are connected to the spin velocity.
In the following, we consider a ferromagnetic system magnetized along the global $z$ direction, and describe the undamped modes for $\alpha=0$. The rotational symmetry around the $z$ axis is preserved in the absence of spin–orbit coupling, and the $z$ component of the total angular momentum, $\sum_{i}L^{z}_{i}$, remains a conserved quantity. On the level of spin-wave theory, this is reflected by the absence of pairing terms in the Hamiltonian $\mathcal{H}_{\textrm{SW}}$, which would couple the first and second components of $\bm{S}^{\bot}$ together. Considering only the $S_{i}^{+}$ sector, two types of excitations may be distinguished in inertial dynamics [47]: precessional excitations performing a counterclockwise rotation around the direction of the total angular momentum, which have $\omega>0$; and nutational excitations performing a clockwise rotation with $\omega<0$. These are illustrated in Fig. 1. Due to the different rotational senses, the angular momenta of the excited states also differ: $\sum_{i}L^{z}_{i}>\sum_{i}\frac{M_{i}}{\gamma}S^{z}_{i}$ for precessional and $\sum_{i}L^{z}_{i}<\sum_{i}\frac{M_{i}}{\gamma}S^{z}_{i}$ for nutational motion. To observe topological edge modes connecting the precessional and nutational magnon bands, this conservation of the total angular momentum has to be broken.
Figure 1: Sketch of the ferromagnet on the honeycomb lattice, with red and blue arrows showing the two sublattices. Red and blue trajectories denote counterclockwise precessional and clockwise nutational modes, respectively.
@@
where $A$ and $B$ indicate the two sublattices denoted by red and blue spins in Fig. 1, respectively. The exchange interaction tensor $\bm{J}_{ij}$ between nearest neighbors is diagonal here, with $J^{zz}<J^{xx}=J^{yy}<0$ preferring magnetization along $\hat{\bm{z}}$. The pseudodipolar interaction $F$ connects the spin directions to the relative positions of the sites via the unit vector $\hat{\bm{r}}_{ij}$ connecting sites $i$ and $j$, which is either $\hat{{\bm{\delta}}}_{1}=\left(-1,0\right)$, $\hat{\bm{\delta}}_{2}=\left(\frac{1}{2},-\frac{\sqrt{3}}{2}\right)$, or $\hat{\bm{\delta}}_{3}=\left(\frac{1}{2},\frac{\sqrt{3}}{2}\right)$. The parameter values used in the simulations are summarized in Table 1. The exchange interaction $J^{xx}=-3$ meV and the magnetic moments $M_{A}=M_{B}=3\mu_{\textrm{B}}$ are comparable to experimentally determined values for honeycomb van der Waals materials [35, 48, 36, 37]. The inertial relaxation time was set to $\eta=330$ fs, which is similar to the values determined experimentally in transition metals [10, 8].
@@
$\mathcal{H}_{\textrm{SW},\bm{k}}=$ $\begin{pmatrix}\Omega&\mathcal{H}_{1}&0&\tilde{F}_{1}-{\rm i}\tilde{F}_{2}\\[10.0pt] \mathcal{H}_{1}^{*}&\Omega&\tilde{F}_{1}^{*}-{\rm i}\tilde{F}^{*}_{2}&0\\[10.0pt] 0&\tilde{F}_{1}+{\rm i}\tilde{F}_{2}&\Omega&\mathcal{H}_{1}\\[10.0pt] \tilde{F}_{1}^{*}+{\rm i}\tilde{F}^{*}_{2}&0&\mathcal{H}_{1}^{*}&\Omega\end{pmatrix}\,$, (5)
where the components are written in the order $\left[S_{-}^{A},S_{-}^{B},S_{+}^{A},S_{+}^{B}\right]$. The matrix elements are given by
@@
$\mathcal{H}_{1}\left(\bm{k}\right)=$ $\left(J^{xx}+\frac{F}{2}\right)\left(\textrm{e}^{\textrm{i}\bm{k}\bm{\delta}_{1}}+\textrm{e}^{\textrm{i}\bm{k}\bm{\delta}_{2}}+\textrm{e}^{\textrm{i}\bm{k}\bm{\delta}_{3}}\right),$ (7)
$\tilde{F}_{1}\left(\bm{k}\right)=$ $\frac{F}{4}\left(2\textrm{e}^{\textrm{i}\bm{k}\bm{\delta}_{1}}-\textrm{e}^{\textrm{i}\bm{k}\bm{\delta}_{2}}-\textrm{e}^{\textrm{i}\bm{k}\bm{\delta}_{3}}\right),$ (8)
$\tilde{F}_{2}\left(\bm{k}\right)=$ $\frac{\sqrt{3}F}{4}\left(\textrm{e}^{\textrm{i}\bm{k}\bm{\delta}_{3}}-\textrm{e}^{\textrm{i}\bm{k}\bm{\delta}_{2}}\right).$ (9)
The calculated magnon spectrum is shown in Fig. 2 along the high-symmetry lines of the Brillouin zone. The exchange interaction forms two precessional bands in this two-sublattice structure, which touch in the $K$ and $K^{\prime}$ points forming Dirac cones. Spin inertia leads to additional nutational modes, which have the same dispersion apart from being shifted up in frequency by $\eta^{-1}$ [17]; see the SM [45]. For sufficiently high values of $\eta\gamma M^{-1}\left|J^{xx}\right|$, first the top of the higher precessional band touches the bottom of the lower nutational band in the $\Gamma$ point, then a nodal line is formed between the bands along a contour encompassing the $\Gamma$ point. The pseudodipolar interaction $F$ breaks the continuous rotational symmetry around $\hat{\bm{z}}$, thereby causing a hybridization between the precessional and nutational modes and opening gaps in the spectrum. In the Dirac points, the gap is opened for any finite value of $F$, but its size scales with $\left|F\right|^{2}$, making it difficult to identify in Fig. 2; see the SM [45]. Along the nodal line, the gap size scales linearly with $\left|F\right|$, but this effect can only be observed if the inertial relaxation time is sufficiently high such that the bands overlap. The hybridization between the nutational and precessional bands may be captured by calculating the ellipticity of mode $n$ at wave vector $\bm{k}$,
@@
Figure 3: Magnon band structure in a slab geometry. The system is $40$ unit cells wide along the $x$ direction with open boundary conditions, while it is infinite along the $y$ direction with periodic boundary conditions, meaning that the eigenstates can be indexed by the wave vector component $k_{y}$. $a$ denotes the distance between nearest-neighbor atoms in the lattice, i.e., the lattice constant is $\sqrt{3}a$. Dotted lines show the boundaries of the one-dimensional Brillouin zone. Coloring shows the localization defined in Eq. (11).
To confirm the topologically non-trivial origin of the edge states, we calculate the topological invariants of the bands in the infinite system, corresponding to the Chern number in the considered case. The Chern number is the integral of the $z$ component of the Berry curvature $C_{n}=\left(2\pi\right)^{-1}\int B^{z}_{n}\left(\bm{k}\right)\textrm{d}^{2}\bm{k}$, where the latter may be expressed with the left and right eigenvectors as (see the SM [45] for details)
@@
Here, $\mathcal{D}_{\bm{k}}$ is the matrix on the right-hand side of Eq. (2) in Fourier space, and $\alpha,\beta,\gamma$ denote Cartesian indices. The distribution of $B^{z}_{n}$ for the four bands is shown in Fig. 4. The sum of the Chern numbers $\left|\sum_{i=1}^{n}C_{i}\right|$ predicts the number of pair of chiral edge modes between bands $n$ and $n+1$, which equals $1$ between any two bands, in agreement with Fig. 3. The Chern number $C_{1}=-1$ of the lowest band in Fig. 4(a) stems from the peaks in the Berry curvature close to the $K$ and $K^{\prime}$ points where the degeneracy is lifted. These peaks are also observable in the second band in Fig. 4(b) with opposite sign, but they are cancelled out by the circular depression around the contour where the precessional and nutational bands overlap in the absence of spin–orbit coupling. The Berry curvatures in the third and fourth bands are qualitatively similar to the second and first bands, respectively, although with reversed signs.
@@
In summary, we demonstrated that the hybridization between nutational and precessional spin waves in inertial spin dynamics may modify the topological properties of the bands. We illustrated this on the example of a honeycomb ferromagnet, which realizes a magnonic Chern insulator with an avoided crossing between the precessional and nutational bands in the bulk system, which is filled by chiral edge modes in a slab geometry.
Note that the width of the magnon bands is lower in the calculations than what was experimentally determined in van der Waals magnets [35, 48, 36, 37]; this is in part caused by the compression of the bands by the inertial term [17], as shown in the SM [45]. This means that the value of the exchange interaction may be higher than what was determined from the non-inertial model, and since the topological transition is governed by the parameter $\gamma M^{-1}\left|J^{xx}\right|\eta$, it may be observed for lower inertial relaxation times than the value used in this work. The inertial relaxation time may be further enhanced in the presence of heavy elements with high spin–orbit coupling in these materials. Therefore, the proposed effects may be experimentally accessible by measuring the magnon spectrum in the whole Brillouin zone up to higher energies where the nutational bands are expected to emerge.
The nutational and precessional bands may be distinguished based on their different angular momentum, or opposite rotational sense. This means that methods sensitive to the rotational sense of the excitations may be particularly suitable for this purpose, such as spin-polarized electron-energy loss spectroscopy. Optical methods may also be applied in the vicinity of the $\Gamma$ point where the reversed rotational sense of the higher modes already provides an indication for the topologically non-trivial nature of the bands. Detecting the inertial dynamics may also shed further light on the role of the competing mechanisms of the Dzyaloshinsky–Moriya interaction and the pseudodipolar interaction in the gap opening in these systems [36, 44]: while the pseudodipolar interaction may open a gap between the precessional and nutational bands, the out-of-plane component of the Dzyaloshinsky–Moriya interaction has no such effect since it preserves the angular momenta of the excitations.
These calculations may be generalized to other systems, for example single-sublattice ferromagnets, where topologically non-trivial bands may only be formed if inertial effects are taken into account; see the SM [45] for an example. The generation of bulk and topological edge modes in ferromagnets at THz frequencies may also open further perspectives in controlling the angular-momentum transfer between magnons and phonons or electrons.